### v1

@@ 5 Simulated Lyman-$\alpha$ datasets as forecast tool
DESI mocks can be used to forecast the constraining power of the DESI Lyman-$\alpha$ forest dataset, particularly for BAO scale measurement uncertainties, in an alternative and mostly complementary way to a Fisher Forecast. Mock forecast have been useful in a number of situations during the DESI survey preparation stages. For instance, to test the gain DESI would obtain if the target selection methods were able to efficiently select fainter quasars than those considered in the DESI nominal design model, with only one pass, or to determine an optimal observing strategy from the perspective of the Lyman-$\alpha$ studies. In this section we perform a mock based forecast of the constraining power of the full DESI survey. We will do so by using the forecast mode of the vega [^18] https://github.com/andreicuceu/vega package [70]. This mode starts with the generation of a simulated noiseless correlation function based on a base model given as input. Such correlation is then paired with the covariance matrix, obtained from equation 4.8, of a mock realization to construct a Gaussian likelihood. Then, for a given correlation function model, with free parameters, to fit this synthetic data, the posterior distribution for the free parameters is sampled using the Polychord [^19] https://github.com/PolyChord/PolyChordLite [71, 72] nested sampler. [^20] Note that in general the model used to simulate the correlation could differ from the model used to fit it. Finally, our forecasted uncertainty for each of the free parameters will be given by the 68% confidence level obtained by analyzing the sampling chains using the GetDist package [73].
@@ 5.1 EDR+M2 mock forecast comparison with measurements
We use the Lyman-$\alpha$ auto and Lyman-$\alpha$ – QSO cross correlations joint fit results obtained by the analysis on observed data (Table 1 in [64]) as our input base model. That is, we use the same correlation function model as in the observed data analysis which includes non-linear BAO broadening, small-scale corrections with the model from [74], contamination due to DLAs and metals among other features (see section 4 of [64] for a full description of the model). The free parameters in this forecast are the bias of Lyman-$\alpha$, HCDs and QSOs ($b_{Ly\alpha}$, $b_{HCD}$ and $b_{QSO}$), the Lyman-$\alpha$ RSD parameter $\beta_{Ly\alpha}$, the velocity bias $b_{\eta,m}$ of each of the metals considered in our mocks, the QSO systematic redshift error shift $\Delta{r}_{\parallel,QSO}$, the statistical redshift errors parameter $\sigma_{v}$, the quasar radiation effect scale $\xi^{TP}_{0}$, the instrumental systematic error amplitude $A_{\rm{inst}}$ and the BAO amplitude $A_{\rm{bao}}$. We use flat priors for all of the free parameters and fit to the correlations functions in the $18010<r<$ 180 Mpc/h range. For the covariance matrix we use one of the LyaCoLoRe mocks realizations which include DLAs, BALs and metals described in section 4. Note that we chose to work with LyaCoLoRe mocks over Saclay ones simply because they cover a higher redshift range. Table 2 presents the parameters that we fit in the model. We show the central value, and the uncertainties of such parameters, that were used to define the base model for the fits, all coming from the measurements in observed data. We also show the uncertainties obtained by our forecast procedure. Although this forecast does not include the BAO scale parameters, $\alpha_{\perp}$ and $\alpha_{\parallel}$, since these were fixed to one in the DESI EDR+M2 analysis due to the relatively low statistical power of observed data [64], we can note that the relative difference between the forecast and measured uncertainties on the free parameters are almost all below the 33% relative difference. We consider that this adequately validates this methodology of using mocks as a forecasting tool, given the differences in the SNR of our mocks compared to data discussed in section 4 which directly contribute on the results of the covariance matrix used to perform the forecast and therefore the obtained uncertainties. Additionally, our mocks have an approximate representation of small-scale clustering producing a smaller variance than data, which could also have a contribution to the differences obtained.
@@ 5.2 Full DESI survey forecast
We use the methodology previously described to simulate Lyman-$\alpha$ quasar spectra across the 14,000 sq.deg expected to be covered by DESI, assuming a fixed exposure time of 4000 seconds for all targets without galactic extinction added to spectra. For direct comparison with the Fisher forecast we use the same throughput model (black line in figure 3), redshift-magnitude distribution and object density as expected by the DESI survey nominal design, hereafter referred as the DESI-Y5 DESIMODEL mock. For completeness we also perform a forecast using the redshift and magnitude distributions as observed in EDR+M2, which we will refer simply as the DESI-Y5 mock. Doing so results in 1.07 million $z>1.8$ QSOs for the DESI-Y5 DESIMODEL mock, and 1.4 million for DESI-Y5. From these, 698k, and 929k, targets are Lyman-$\alpha$ ($z>2.1$) for each mock respectively. Figure 15 shows the resulting number density of quasars simulated over the DESI footprint, as expected by the observed EDR+M2 redshift and magnitude distributions divided into nside=16 HEALpix pixels.
@@
We opted to generate mocks including only Lyman-$\alpha$ absorptions in order to compare to the results of the Fisher forecast formalism presented in [28], which did not include any sort of contaminants in its assumptions. The same forecasts could be performed including contaminants, however this requires further study on the effect of the involved parameter values set on the base model on the resulting uncertainties and goes beyond the scope of this paper.
@@
In figure 16 we show the forecasted uncertainty through the 68% and 95% confidence level contour plots of the BAO parameters $\alpha_{\perp}$ and $\alpha_{\parallel}$ for both models and compare with the measurements of DR16 [10] to highlight the expected constraining power of the completed DESI survey. The results of our DESI-Y5 mock forecasts have a relative difference with the Fisher forecast performed in  [28] of a 17% for $\nicefrac{{\sigma_{Hr_{d}}}}{{Hr_{d}}}$ and 8% for $\nicefrac{{\sigma_{\nicefrac{{D_{A}}}{{r_{d}}}}}}{{\nicefrac{{D_{A}}}{{r_{d}}}}}$. Regarding the DESI-Y5 mock the results have a relative difference with respect of the Fisher forecast performed in [29] of 28% for $\nicefrac{{\sigma_{Hr_{d}}}}{{Hr_{d}}}$ and 20% for $\nicefrac{{\sigma_{\nicefrac{{D_{A}}}{{r_{d}}}}}}{{\nicefrac{{D_{A}}}{{r_{d}}}}}$.
@@ 6 Summary and Conclusions
We have presented our methodology to produce synthetic DESI Lyman-$\alpha$ datasets. This methodology implements the use of the script quickquasars, which compiles multiple modules from the desisim and specsim repositories of the DESI main code dedicated to produce synthetic Lyman-$\alpha$ spectra, and has the following characteristics:
@@
It is important to mention that mocks produced with quickquasars have been used to test systematics in the analysis pipelines and as validation steps on other works before. The eBOSS DR16 [10] BAO study with Lyman-$\alpha$ forest included the use of eBOSS mocks as a validation step, these same mocks were used to test the efficiency and purity of the eBOSS DR16 DLA catalog [52]. DESI mocks have been used to study the effect of quasar redshift errors on Lyman-$\alpha$ forest correlation functions [37], test the efficiency and purity of the BAL [48] and DLA [53] finder algorithms, study the impact of BALs on quasar redshift measurements [78], and for validation of P1D studies [13, 12].
@@ Acknowledgments
We thank Ian McGreer for his valuable contributions to integrate simqso into the desisim package and consequently into quickquasars.
@@
This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Science and Technology of Mexico (CONACYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies.
@@
Data Points
The data points corresponding to each figure in this paper can be accessed in the Zenodo repository at https://doi.org/10.5281/zenodo.10433340
@@ Appendix A Emission Lines included in simulated spectra
For the EDR+M2 and DESI-Y5 mocks produced in this work we have defined the emission lines as the combination of two emission line models. One the one hand, we use emission lines model that are within the Lyman-$\alpha$ region of the composite model of the BOSS spectra [43] which was computed from over 102k QSOs in the redshift range of $2.1\leq z\leq 3.5$ and a rest-frame wavelength range of $800\text{\AA}<\lambda<3300\text{\AA}$, not including BAL or DLA quasars. On the other hand, outside the Lyman-$\alpha$ region we use the version 7 emission lines model of simqso which includes several datasets of QSO spectra observations on a wider rest-frame wavelength range than the BOSS spectra composite model. We refer the reader to the main simqso repository (given in footnote 9) for further details on the datasets used to construct this model. The resulting model used for our mocks is shown in table A.1. We highlight those lines that correspond to the composite model of [43] within the Lyman-$\alpha$ region. The equivalent width of some lines were tuned to resemble the composite model obtained using EDR+M2 spectra. Figure A.1 displays a qualitative comparison between composite models of DESI EDR+M2 observed data and a mock realization whose details are given in section 4. We can note that mocks contain the main features of QSO continuum.
@@ Appendix B Raw transmission data model
In order for quickquasars to work the input, FITS transmission files should follow the data model of the Raw Mocks. These files are usually divided by HEALpix pixel following the name convention of transmission-<nside>-<healpix>.fits.gz, where <nside> is the used nside of HEALpix pixels (nside=16 for the LyaCoLoRe, Saclay and Ohio raw mocks) in NESTED scheme and <healpix> is the index of the HEALpix.
@@
METADATA [^22] This HDU must include the following information in its header: HPXPIXEL (int): HEALpix pixel index of the transmission file. HPXNEST (boolean): True if the NESTED scheme of HEALpix was used. HPXNSIDE (int): HEALpix nside used for creating the transmission files.
BinTableHDU
Metadata of the mock QSOs (see table B.2).
WAVELENGTH
ImageHDU
Wavelength $\lambda$ of the transmitted flux in angstroms.
F_LYA
ImageHDU
Lyman-$\alpha$ absorptions transmitted flux. Can also be called TRANSMISSION.
F_LYB
ImageHDU
Lyman-$\beta$ absorptions transmitted flux. Optional.
F_METALS
ImageHDU
Metal absorptions transmitted flux. Optional.
DLA
BinTableHDU
Metadata of the correlated DLAs (see table B.2). Optional.
@@
RA
FLOAT
Right Ascension in degrees.
DEC
FLOAT
Declination in degrees.
Z_noRSD
FLOAT
Redshift without RSD effects.
Z
FLOAT
Redshift.
MOCKID
INT
Mock QSO identification number.
@@
Z_DLA_NO_RSD
DOUBLE
DLA redshift without RSD effects.
Z_DLA_RSD
DOUBLE
DLA redshift.
N_HI_DLA
DOUBLE
DLA hydrogen column density.
MOCKID
INT
Mock quasar identification number.
DLAID
INT
Mock DLA identification number.
@@ Appendix C Quickquasars output files data model
The truth FITS files follow the data format given in table C.1, the files contain information about all the generated quasars, except for the DLA and BAL HDUs which only include quasars including these features. Many of the entries in these files are added only if the corresponding feature was used, i.e BALs and DLAs or an option to store the information is given, i.e continuum templates and resolution matrices. Table C.2 contains the format of each of the BinTableHDU entries of the truth files.
@@
TRUTH
BinTableHDU
Truth values of the generated quasars.
TRUTH_QSO
BinTableHDU
Quasars supplemental metadata.
DLA_META*
BinTableHDU
DLA metadata.
BAL_META*
BinTableHDU
BAL templates metadata.
TRUE_CONT*
BinTableHDU
Continuum templates information.
B_RESOLUTION*
ImageHDU
B-band resolution matrix.
R_RESOLUTION*
ImageHDU
R-band resolution matrix.
Z_RESOLUTION*
ImageHDU
Z-band resolution matrix.
@@
TARGETID
INT
Target identification number.
OBJTYPE
STR
Spectral type. QSO by default.
SUBTYPE
STR
Spectral sub-type (e.g LYA, LRG, ELG). Left blank by default.
TEMPLATEID
INT
Continuum template identification number.
SEED
FLOAT
Random seed used to generate spectrum.
Z
DOUBLE
Redshift.
MAG
FLOAT
Spectrum normalization magnitude.
MAGFILTER
STR
Spectrum normalization filter.
FLUX_G
FLOAT
Flux in G-band.
FLUX_R
FLOAT
Flux in R-band.
FLUX_Z
FLOAT
Flux in Z-band.
FLUX_W1
FLOAT
Flux in WISE W1-band.
FLUX_W2
FLOAT
Flux in WISE W2-band.
TRUEZ
FLOAT
Redshift without shifts due to statistical imprecision or systematic shifts.
Z_INPUT
DOUBLE
Redshift without Fingers of God effect.
DZ_FOG
DOUBLE
Redshift shift due to Finger of God effect.
DZ_SYS
DOUBLE
Systematic shift applied to redshift. 0 by default.
DZ_STAT
DOUBLE
Redshift shift emulating the statistical imprecision of redrock. Present only if applied.
Z_NORSD
FLOAT
Like Z_INPUT but without RSD effects.
EXPTIME
FLOAT
Exposure time. Present only if the exposure time was assigned using a distribution.
@@
TARGETID
INT
Target identification number.
BAL_TEMPLATEID
INT
BAL template identification number. Set to -1 if the quasar is not a BAL.
DLA
BOOLEAN
True if the quasar hosts a DLA.
@@
MABS_1450
FLOAT
Rest-frame absolute magnitude at 1450 angstroms.
SLOPES
FLOAT [5]
Broken power law continuum slopes.
EMLINES
FLOAT [3,M [^27] Denotes the number of emission lines present in the emission line model (73 in our case, see table A.1).]
Emission line parameters used to generate the continuum.
@@
PCA_COEFF
FLOAT [4]
PCA coefficients used to generate the continuum.
@@
NHI
DOUBLE
DLA hydrogen column density.
Z_DLA
DOUBLE
DLA redshift.
TARGETID
INT
Target identification number.
DLAID
INT
DLA identification number.
@@
TARGETID
INT
Target identification number.
Z
FLOAT
Redshift.
BAL_PROB
FLOAT
BAL probability. Set to 1 by default.
BAL_TEMPLATEID
INT
BAL template identification number.
BI_CIV
FLOAT
C IV Balnicity Index.
ERR_BI_CIV
FLOAT
C IV Balnicity Index uncertainty.
NCIV_2000
INT
Number of troughs wider that 2000 km/s.
VMIN_CIV_2000
FLOAT [5]
Minimum velocity of each absorption trough.
VMAX_CIV_2000
FLOAT [5]
Maximum velocity of each absorption trough.
POSMIN_CIV_2000
FLOAT [5]
Position of the minimum of each absorption trough.
FMIN_CIV_2000
FLOAT [5]
Normalized flux density at the minimum of each absorption trough.
AI_CIV
FLOAT
Absorption Index.
ERR_AI_CIV
FLOAT
Absorption Index uncertainty.
NCIV_450
INT
Number of troughs wider that 450 km/s.
VMIN_CIV_450
FLOAT [27]
Minimum velocity of each absorption trough.
VMAX_CIV_450
FLOAT [27]
Maximum velocity of each absorption trough.
POSMIN_CIV_450
FLOAT [27]
Position of the minimum of each absorption trough.
FMIN_CIV_450
FLOAT [27]
Normalized flux density at the minimum of each absorption trough.
@@
TARGETID
INT
Target identification number.
TRUE_CONT
DOUBLE [W [^29] Denotes the wavelength array size; it should be the same as the output wavelength arrays of quickquasars. Default to 3251 for DESI mocks.]
Continuum template.

### v3

@@ 5 Simulated Lyman-$\alpha$ datasets as forecast tool
DESI mocks can be used to forecast the constraining power of the DESI Lyman-$\alpha$ forest dataset, particularly for BAO scale measurement uncertainties, in an alternative and mostly complementary way to a Fisher Forecast. Mock forecast have been useful in a number of situations during the DESI survey preparation stages. For instance, to test the gain DESI would obtain if the target selection methods were able to efficiently select fainter quasars than those considered in the DESI nominal design model, with only one pass, or to determine an optimal observing strategy from the perspective of the Lyman-$\alpha$ studies. In this section, we perform a mock based forecast of the constraining power of the full DESI survey. We will do so by using the forecast mode of the vega [^18] https://github.com/andreicuceu/vega package [82]. This mode starts with the generation of a simulated noiseless correlation function based on a base model given as input. The simulated correlation is then paired with the covariance matrix, obtained from equation 4.8, of a mock realization to construct a Gaussian likelihood. Then, for a given correlation function model, with free parameters, to fit this synthetic data, the posterior distribution for the free parameters is sampled using the Polychord [^19] https://github.com/PolyChord/PolyChordLite [83, 84] nested sampler. [^20] Note that in general the model used to simulate the correlation could differ from the model used to fit it. Finally, our forecasted uncertainty for each of the free parameters will be given by the 68% credible region obtained by analyzing the sampling chains using the GetDist package [85].
@@ 5.1 EDR+M2 mock forecast comparison with measurements
We use the Lyman-$\alpha$ auto and Lyman-$\alpha$ – QSO cross correlations joint fit results obtained by the analysis on observed data (Table 1 in [76]) as our input base model. That is, we use the same correlation function model as in the observed data analysis which includes non-linear BAO broadening, small-scale corrections with the model from [86], contamination due to DLAs and metals among other features (see section 4 of [76] for a full description of the model). The free parameters in this forecast are the bias of Lyman-$\alpha$, HCDs and QSOs ($b_{Ly\alpha}$, $b_{HCD}$ and $b_{QSO}$), the Lyman-$\alpha$ RSD parameter $\beta_{Ly\alpha}$, the velocity bias $b_{\eta,m}$ of each of the metals considered in our mocks, the QSO systematic redshift error shift $\Delta{r}_{\parallel,QSO}$, the statistical redshift errors parameter $\sigma_{v}$, the quasar radiation effect scale $\xi^{TP}_{0}$, the instrumental systematic error amplitude $A_{\rm{inst}}$ and the BAO amplitude $A_{\rm{bao}}$. We use flat priors for all free parameters and fit the correlation functions in the $18010<r<$ 180 Mpc/h range. For the covariance matrix, we use one of the LyaCoLoRe mocks realizations which include DLAs, BALs and metals described in section 4. Note that we chose to work with LyaCoLoRe mocks over Saclay ones simply because they cover a higher redshift range. Table 2 presents the free parameters in the model whose posterior distribution is sampled. We show the central values, that were used to define the base model to perform our forecast, and the uncertainties of such parameters all coming from the measurements in observed data. We also show the uncertainties obtained by our forecast procedure. Although this forecast does not include the BAO scale parameters, $\alpha_{\perp}$ and $\alpha_{\parallel}$, since these were fixed to one in the DESI EDR+M2 analysis due to the relatively low statistical power of the data sample [76], we can note that the relative difference between the uncertainties obtained by our forecast and the observed data case, defined as $\Delta_{\sigma}=|1-\sigma_{\rm{Forecast}}/\sigma_{\rm{obs}}|$, are almost all below the 33% level. We consider that this adequately validates this methodology of using mocks as a forecasting tool, given the differences in the SNR of our mocks compared to data discussed in section 4 which directly contribute on the results of the covariance matrix used to perform the forecast and therefore the obtained uncertainties. Additionally, our mocks have an approximate representation of small-scale clustering producing a smaller variance than data, which could also have a contribution to the differences obtained.
@@ 5.2 Full DESI survey forecast
We use the methodology previously described to simulate Lyman-$\alpha$ quasar spectra across the 14,000 sq.deg expected to be covered by DESI, assuming a fixed exposure time of 4000 seconds for all targets without galactic extinction added to spectra. For direct comparison with the Fisher forecast we use the same throughput model (black line in figure 3), redshift-magnitude distribution and object density as expected by the DESI survey nominal design, hereafter referred as the DESI-Y5 DESIMODEL mock. For completeness we also perform a forecast using the redshift and magnitude distributions as observed in EDR+M2, which we will refer simply as the DESI-Y5 mock. Doing so results in 1.07 million $z>1.8$ QSOs for the DESI-Y5 DESIMODEL mock, and 1.4 million for DESI-Y5. From these, 698k, and 929k, targets are Lyman-$\alpha$ ($z>2.1$) for each mock, respectively. Figure 15 shows the resulting number density of quasars simulated over the DESI footprint, as expected by the observed EDR+M2 redshift and magnitude distributions divided into nside=16 HEALpix pixels.
@@
We opted to generate mocks including only Lyman-$\alpha$ absorptions in order to compare to the results of the Fisher forecast formalism presented in [42], which did not include any sort of contaminants in its assumptions. The same forecasts could be performed including contaminants, however this requires further study on the effect of the involved parameter values set on the base model on the resulting uncertainties and goes beyond the scope of this manuscript.
@@
In figure 16 we show the forecasted uncertainty through the 68% and 95% credible regions of the BAO parameters $\alpha_{\perp}$ and $\alpha_{\parallel}$ for both models and compare with the measurements of DR16 [10] to highlight the expected constraining power of the completed DESI survey. The results of our DESI-Y5 mock forecasts have a relative difference with the Fisher forecast performed in  [42] of a 17% for $\nicefrac{{\sigma_{Hr_{d}}}}{{Hr_{d}}}$ and 8% for $\nicefrac{{\sigma_{\nicefrac{{D_{A}}}{{r_{d}}}}}}{{\nicefrac{{D_{A}}}{{r_{d}}}}}$. Regarding the DESI-Y5 mock the results have a relative difference with respect of the Fisher forecast performed in [44] of 28% for $\nicefrac{{\sigma_{Hr_{d}}}}{{Hr_{d}}}$ and 20% for $\nicefrac{{\sigma_{\nicefrac{{D_{A}}}{{r_{d}}}}}}{{\nicefrac{{D_{A}}}{{r_{d}}}}}$.
@@ 6 Summary and Conclusions
We have presented the methodology to produce synthetic DESI Lyman-$\alpha$ datasets. This methodology implements the use of the script quickquasars, which compiles multiple modules from the desisim and specsim repositories of the DESI main code dedicated to produce synthetic Lyman-$\alpha$ spectra, and has the following characteristics:
@@
It is important to mention that mocks produced with quickquasars have been used to test systematics in the analysis pipelines and as validation steps in other works before. The eBOSS DR16 [10] BAO study with Lyman-$\alpha$ forest included the use of eBOSS mocks as a validation step, these same mocks were used to test the efficiency and purity of the eBOSS DR16 DLA catalog [64]. DESI mocks have been used to study the effect of quasar redshift errors on Lyman-$\alpha$ forest correlation functions [49], test the efficiency and purity of the BAL [60] and DLA [65] finder algorithms, study the impact of BALs on quasar redshift measurements [90], and for validation of P1D studies [20, 19].
@@
Data Points
The data points corresponding to each figure in this paper can be accessed in the Zenodo repository at https://doi.org/10.5281/zenodo.10433340.
@@ Acknowledgments
The authors thank Ian McGreer for his contributions and efforts to integrate simqso into the desisim package and consequently into quickquasars. Special thanks to James Rich and Julian Bautista for their invaluable comments and discussions on this manuscript. The authors would also like to thank James Farr, Hélion du Mas des Bourboux, Thomas Etourneau, and all the persons involved on the production of the LyaCoLoRe and Saclay mocks.
@@
This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies.
@@ Appendix A Emission Lines included in simulated spectra
For the EDR+M2 and DESI-Y5 mocks produced in this work we have defined the emission lines as the combination of two emission line models. One the one hand, we use emission lines model that are within the Lyman-$\alpha$ region of the composite model of the BOSS spectra [55] which was computed from over 102k QSOs in the redshift range of $2.1\leq z\leq 3.5$ and a rest-frame wavelength range of $800\text{\AA}<\lambda<3300\text{\AA}$, not including BAL or DLA quasars. On the other hand, outside the Lyman-$\alpha$ region we use the version 7 emission lines model of simqso which includes several datasets of QSO spectra observations on a wider rest-frame wavelength range than the BOSS spectra composite model. We refer the reader to the main simqso repository (given in footnote 9) for further details on the datasets used to construct this model. The resulting model used for our mocks is shown in LABEL:tab:emlinemodel. We highlight those lines that correspond to the composite model of [55] within the Lyman-$\alpha$ region. The equivalent width of some lines were tuned to resemble the composite model obtained using EDR+M2 spectra. Figure A.1 displays a qualitative comparison between composite models of DESI EDR+M2 observed data and a mock realization whose details are given in section 4. We can note that mocks contain the main features of QSO continuum.
@@ Appendix B Raw transmission data model
In order for quickquasars to work, the input FITS transmission files should follow the data model of the raw mocks given in Table B.1. These files are usually divided by HEALpix pixel following the name convention of transmission-<nside>-<healpix>.fits.gz, where <nside> is the used nside of HEALpix pixels (nside=16 for the LyaCoLoRe, Saclay and Ohio raw mocks) in NESTED scheme and <healpix> is the index of the HEALpix.
@@
METADATA [^22] This HDU must include the following information in its header: HPXPIXEL (int): HEALpix pixel index of the transmission file. HPXNEST (boolean): True if the NESTED scheme of HEALpix was used. HPXNSIDE (int): HEALpix nside used for creating the transmission files. BinTableHDU Metadata of the mock QSOs (see table B.2).
WAVELENGTH ImageHDU Wavelength $\lambda$ of the transmitted flux in angstroms.
F_LYA ImageHDU Lyman-$\alpha$ absorptions transmitted flux. Can also be called TRANSMISSION.
F_LYB ImageHDU Lyman-$\beta$ absorptions transmitted flux. Optional.
F_METALS ImageHDU Metal absorptions transmitted flux. Optional.
DLA BinTableHDU Metadata of the correlated DLAs (see table B.2). Optional.
@@
RA FLOAT Right Ascension in degrees.
DEC FLOAT Declination in degrees.
Z_noRSD FLOAT Redshift without RSD effects.
Z FLOAT Redshift.
MOCKID INT Mock QSO identification number.
@@
Z_DLA_NO_RSD DOUBLE DLA redshift without RSD effects.
Z_DLA_RSD DOUBLE DLA redshift.
N_HI_DLA DOUBLE DLA hydrogen column density.
MOCKID INT Mock quasar identification number.
DLAID INT Mock DLA identification number.
@@ Appendix C Quickquasars output files data model
The truth FITS files follow the data format given in table C.1, the files contain information about all the generated quasars, except for the DLA and BAL HDUs which only include quasars including these features. Many of the entries in these files are added only if the corresponding feature was used, i.e BALs and DLAs or an option to store the information is given, i.e continuum templates and resolution matrices. LABEL:tab:truthformat contains the format of each of the BinTableHDU entries of the truth files.
@@
TRUTH BinTableHDU Truth values of the generated quasars.
TRUTH_QSO BinTableHDU Quasars supplemental metadata.
DLA_META* BinTableHDU DLA metadata.
BAL_META* BinTableHDU BAL templates metadata.
TRUE_CONT* BinTableHDU Continuum templates information.
B_RESOLUTION* ImageHDU B-band resolution matrix.
R_RESOLUTION* ImageHDU R-band resolution matrix.
Z_RESOLUTION* ImageHDU Z-band resolution matrix.
@@
TARGETID INT Target identification number.
OBJTYPE STR Spectral type. QSO by default.
SUBTYPE STR Spectral sub-type (e.g LYA, LRG, ELG). Left blank by default.
TEMPLATEID INT Continuum template identification number.
SEED FLOAT Random seed used to generate spectrum.
Z DOUBLE Redshift.
MAG FLOAT Spectrum normalization magnitude.
MAGFILTER STR Spectrum normalization filter.
FLUX_G FLOAT Flux in G-band.
FLUX_R FLOAT Flux in R-band.
FLUX_Z FLOAT Flux in Z-band.
FLUX_W1 FLOAT Flux in WISE W1-band.
FLUX_W2 FLOAT Flux in WISE W2-band.
TRUEZ FLOAT Redshift without shifts due to statistical imprecision or systematic shifts.
Z_INPUT DOUBLE Redshift without Fingers of God effect.
DZ_FOG DOUBLE Redshift shift due to Finger of God effect.
DZ_SYS DOUBLE Systematic shift applied to redshift. 0 by default.
DZ_STAT DOUBLE Redshift shift emulating the statistical imprecision of redrock. Present only if applied.
Z_NORSD FLOAT Like Z_INPUT but without RSD effects.
EXPTIME FLOAT Exposure time. Present only if the exposure time was assigned using a distribution.
@@
TARGETID INT Target identification number.
BAL_TEMPLATEID INT BAL template identification number. Set to -1 if the quasar is not a BAL.
DLA BOOLEAN True if the quasar hosts a DLA.
@@
MABS_1450 FLOAT Rest-frame absolute magnitude at 1450 angstroms.
SLOPES FLOAT [5] Broken power law continuum slopes.
EMLINES FLOAT [3,M [^27] Denotes the number of emission lines present in the emission line model (73 in our case, see LABEL:tab:emlinemodel).] Emission line parameters used to generate the continuum.
@@
PCA_COEFF FLOAT [4] PCA coefficients used to generate the continuum.
@@
NHI DOUBLE DLA hydrogen column density.
Z_DLA DOUBLE DLA redshift.
TARGETID INT Target identification number.
DLAID INT DLA identification number.
@@
TARGETID INT Target identification number.
Z FLOAT Redshift.
BAL_PROB FLOAT BAL probability. Set to 1 by default.
BAL_TEMPLATEID INT BAL template identification number.
BI_CIV FLOAT C IV Balnicity Index.
ERR_BI_CIV FLOAT C IV Balnicity Index uncertainty.
NCIV_2000 INT Number of troughs wider that 2000 km/s.
VMIN_CIV_2000 FLOAT [5] Minimum velocity of each absorption trough.
VMAX_CIV_2000 FLOAT [5] Maximum velocity of each absorption trough.
POSMIN_CIV_2000 FLOAT [5] Position of the minimum of each absorption trough.
FMIN_CIV_2000 FLOAT [5] Normalized flux density at the minimum of each absorption trough.
AI_CIV FLOAT Absorption Index.
ERR_AI_CIV FLOAT Absorption Index uncertainty.
NCIV_450 INT Number of troughs wider that 450 km/s.
VMIN_CIV_450 FLOAT [27] Minimum velocity of each absorption trough.
VMAX_CIV_450 FLOAT [27] Maximum velocity of each absorption trough.
POSMIN_CIV_450 FLOAT [27] Position of the minimum of each absorption trough.
FMIN_CIV_450 FLOAT [27] Normalized flux density at the minimum of each absorption trough.
@@
TARGETID INT Target identification number.
TRUE_CONT DOUBLE [W [^29] Denotes the wavelength array size; it should be the same as the output wavelength arrays of quickquasars. Default to 3251 for DESI mocks.] Continuum template.